Financial Management:

The equity multiplier measures:

operating efficiency.

returns to stockholders.

management efficiency.

financial leverage.

asset use efficiency.
Answer #1

Equity multiplier is found out by dividing total assets of a company by total shareholder's equity. It gives us the information about how the assets of an organization have been financed. It gives us idea about the risk of company and helps in estimating leverage of a company from point of view of creditors or measuring financial leverage

Debt ratio = Debt / Total assets =1 - (Equity/Total assets) = 1 - [1/(total Assets/Equity)] = 1 - (1/Equity multiplier)

As can been from above relation higher equity multiplier will increase the debt ratio and increase financial leverage. On the other hand lower equity multiplier will decrease debt ratio and decrease financial leverage

Hence Equity multiplier measures financial leverage

Answer: Financial leverage